If a scientist has 35% acid solution and 20% acid solution and she needs 3 liters of 25% acid solution write and solve a linear equation and how much of each solution she should use.
Answer by stanbon(75887)

If a scientist has 35% acid solution and 20% acid solution and she needs 3 liters of 25% acid solution write and solve a linear equation and how much of each solution she should use.
----------
Equation:
acid + acid = acis
0.35x + 0.20(3-x) = 0.25*3
-------
35x + 20*3 - 20x = 25*3
15x = 5*3
x = 1 liter (amt. of 35% solution needed)
-------------------------------
3-x = 2 liters (amt. of 20% solution needed)
================================
Cheers,
Stan H.